### 20 Mar 2026 v1

# spacing=nonfrench

Abstract
Formal verification via interactive theorem proving is increasingly used to ensure the correctness of critical systems, yet constructing large proof scripts remains highly manual and limits scalability. Advances in large language models (LLMs), especially in mathematical reasoning, make their integration into software verification increasingly promising. This paper introduces a neuro-symbolic proof generation framework designed to automate proof search for systems-level verification projects. The framework performs a best-first tree search over proof states, repeatedly querying an LLM for the next candidate proof step. On the neural side, we fine-tune LLMs using datasets of proof state–step pairs; on the symbolic side, we incorporate a range of ITP tools to repair rejected steps, filter and rank proof states, and automatically discharge subgoals when search progress stalls. This synergy enables data-efficient LLM adaptation and semantics-informed pruning of the search space. We implement the framework on a new Isabelle REPL that exposes fine-grained proof states and automation tools, and evaluate it on the FVEL seL4 benchmark and additional Isabelle developments. On seL4, the system proves up to 77.6% of the theorems, substantially surpassing previous LLM-based approaches and standalone Sledgehammer, while solving significantly more multi-step proofs. Results across further benchmarks demonstrate strong generalization, indicating a viable path toward scalable automated software verification.
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding authors.
1 Introduction
Formal verification plays a vital role in ensuring the correctness and reliability of software systems, especially in safety- and security-critical domains where failures can be catastrophic (Clarke1996, ; Blanchet2003, ; Woodcock2009, ). Interactive theorem proving (ITP) stands out for its expressiveness and ability to provide the strongest assurance. With ITPs, landmark projects such as CompCert (a formally verified optimizing C compiler whose correctness is mechanically proven in Rocq leroy2016compcert ) and seL4 (formally verified OS microkernel whose functional correctness, security properties and reliability guarantees are proven in Isabelle/HOL klein2014comprehensive ) have been successfully delivered, which demonstrate that formally verified systems can achieve both practical usability and rigorous correctness.
While ITP offers unparalleled assurance, it also incurs substantial costs, preventing its wide adoption, especially in large-scale, systems-level software (ringer2019qed, ). Specifically, a typical software verification workflow involves 1) writing precise formal specifications to be proved and 2) constructing rigorous proofs for these specifications, both of which demand immense human effort. For example, the seL4 verification project required approximately 20 person-years of work, and produced over 100K lines of proof scripts, compared to only about 10K lines of C implementation and merely 3K lines in its simplest abstract specification in Isabelle klein2014comprehensive . Moreover, developing such specifications and proofs necessitates specialized expertise in both theorem proving and the target domain, a combination that is challenging to acquire.
In this work, we focus on the proof generation step in systems-level software verification projects. Recently, large language models (LLMs) have demonstrated encouraging mathematical reasoning capabilities, which inspire research into their application for software verification, particularly for proof generation  qin2025can ; thompson2024rango ; zhang2024selene ; lin2024fvel . Unfortunately, both prior studies and our initial experiments show that, even when theorem statements are well formalized and relevant libraries are fully provided in advance, current LLMs still struggle to generate complete, correct proofs. We summarize two challenges as follows.
❶ Lack of specialized expertise. The first challenge lies in that software verification often relies on numerous domain-specific lemmas and specialized proof tactics. Although modern LLMs perform well in general reasoning tasks, they are not well-versed in these specialized domains, leading to limited effectiveness when applied to proving corresponding theorems. For example, LLM4FSCQ qin2025can evaluated models such as GPT-4o and Gemini 1.5 Flash on FSCQ (a Rocq-based file system verification project fscq ) and achieved only 38% proof coverage. Moreover, in the larger seL4 project, which involves extensive lemma libraries and customized tactics, Selene zhang2024selene found that GPT-4 succeeded on only 20% of a benchmark theorem set, despite the fact that nearly 40% of them require one-line proofs.
❷ Pausity of usable data. The second challenge, which is closely related to the first, is the shortage of high-quality (training) data for software verification. For instance, the seL4 project contains only around 20K theorems and lemmas, and most proofs are written in a procedural style, where much of the reasoning is implicit and encoded in the interaction with the theorem prover. This substantially limits their usefulness in effective LLM training/fine-tuning. Indeed, FVEL lin2024fvel , which systematically extracted theorems from seL4 to fine-tune Mistral-7B-Instruct jiang2023mistral and Llama-3-8B-Instruct dubey2024llama models, reported a success rate below 10% (88 out of 1,077 theorems in the test set). [^1] Selene only extracts a small subset of the seL4 project, and its collected theorems are relatively easy, with proof lengths ranging from one to five. In contrast, the FVEL benchmark is more comprehensive and challenging, as it extracts all the theorems to be proved from seL4.
Approach. In this paper, we propose a proof generation framework that synergistically combines LLMs and ITPs, aiming for automated systems-level software verification. Our approach adopts a proof step-based tree-search pipeline that iteratively queries a fine-tuned LLM to predict the next proof step until the target theorem is successfully proved.
To address the first challenge, our framework adopts a neuro-symbolic approach tailored to the nature of software verification. First, to address challenges posed by domain-specific lemmas and tactics, we incorporate a proof-step revision module that repairs or refines LLM-generated steps whenever they fail to compile or make no progress. Second, to efficiently prune unpromising branches, we utilize Nitpick (blanchette2010nitpick, ) and QuickCheck (claessen2000quickcheck, ) to “test” each candidate proof state and discard those that admit counterexamples; the remaining candidates are then ranked by the LLM according to their plausibility. Finally, if the tree search still fails to resolve the goal, Sledgehammer (bohme2010sledgehammer, ) acts as a backstop that scans the local library for potentially relevant lemmas and attempts to complete the current proof.
To address the second challenge, our tree-search pipeline facilitates direct interaction between the LLM and the theorem prover, by extracting the prover’s internal proof state to LLM, which can inform and guide subsequent step predictions. This pipeline also enables “bootstrapping” training data generation. Namely, during the tree-search process, training data for proof-step prediction can be extracted on-the-fly. For example, if a theorem’s proof consists of five steps, a successful tree search could yield five training instances, which can later be used for, e.g., fine-tuning (wu2024internlm2, ; xin2025bfs, ).
Experiments. We evaluate our framework on seL4 klein2014comprehensive proof construction. The results show that, with a fine-tuned language model, our approach achieves a total proof success rate of up to 77.6%, significantly surpassing prior automated proof-generation approaches for seL4. Noteably, it excels, compared to the two preceding systems, at successfully handling a larger number of multi-step proofs, a challenge that has not been addressed by prior work. Moreover, our framework substantially alleviates manual proof-engineering effort, reducing expert proof-writing effort by 71.1% on average in an AI–human collaboration setting. Finally, experimental results on benchmarks from Archive of Formal Proofs projects (e.g., X86 Semantics , IEEE Floating Point, and SATSolverVerification), as well as the code verification benchmark Code2Inv (by translating its loop invariants into Isabelle theorems) confirm the strong generalizability of our approach.
Summary of main contributions. (1) We provide a neuro-symbolic framework based on LLM-enabled tree search, tailored for real-world, systems-level software verification projects. (2) We implement concrete integration of LLMs and a broad range of symbolic, logic-based tools, substantially enhancing the efficacy of proof-tree search; (3) We curate training datasets and benchmarks, and carry out thorough empirical evaluations. Our work advances a new frontier of AI4Verification, an emerging area focused on applying cutting-edge AI techniques to hardware/software verification, especially to support large-scale proof engineering for safety-critical systems and architectures.
Figure 1: An example theorem and its human-written proof from the seL4 project, shown alongside LLM-generated attempts. The human-written Isabelle proof illustrates the procedural, tactic-driven reasoning style that is common in software verification. We prompted two state-of-the-art LLMs, Gemini 3 and GPT-5.1, to synthesize the same proof under controlled conditions, with web search disabled to avoid any leakage of the original script. Although both models correctly recognized the proof context and invoked some relevant tactics, neither succeeded in producing a valid proof even after roughly one minute of internal reasoning. Their attempts frequently misuse the domain-specific wp tactic, or hallucinate nonexistent lemmas.
Structure. Section 2 illustrates the challenges of the problem. Section 3 presents our neuro-symbolic approach. Section 4 reports the evaluation of our approach on seL4 and four additional repository-level benchmarks. Section 5 reviews the related work. Section 6 discusses the limitations of the current work and outlines directions for future research. Section 7 concludes the paper.
Figure 2: Overview of our neuro-symbolic proof-search framework. Starting from the initial proof state, the system repeatedly generates candidate proof steps using a fine-tuned LLM, followed by rule-based symbolic revision to repair syntactic or semantic issues in the proposed steps. Each accepted step is executed by Isabelle/HOL to derive new proof states, which are then filtered by symbolic tools to prune states containing counterexamples or duplicates. The remaining states are ranked using cumulative LLM log-probabilities, and the highest-scoring ones are selected for further expansion in a best-first search. If the search cannot complete the proof, the framework finally invokes Sledgehammer to attempt to resolve the remaining goals.
2 Problem
To show why existing LLMs struggle to synthesize proofs in real-world software verification tasks, we present a representative theorem from seL4. We choose seL4 as it remains the most comprehensive and influential milestone for system software verification. Its entire C implementation is proven functionally correct against a formal specification in Isabelle/HOL, ensuring the absence of crashes and undefined behavior. Building on this foundation, seL4’s verification was extended to enforce noninterference of information flows and access control, establishing formal guarantees of confidentiality and integrity. To date, seL4 remains one of the safest OS kernels with machine-checked proofs of both implementation-level correctness and security properties, setting the benchmark for formal system verification.
An example from seL4’s verification is demonstrated in Figure 1, where the theorem cap_swap_valid_objs is drawn from the invariant proofs of the abstract specification of the seL4 microkernel, and similar proof patterns recur throughout the project. To prove this theorem, human experts write down (in Isabelle) three proof steps, each comprising one tactic (e.g., simp and wp) and some used lemmas (e.g., cap_swap_def and set_cap_valid_objs; cf. the left panel in Figure 1).
Unlike formal mathematics, theorems in software verification are often formulated as Hoare triples {P} C {Q}, where P and Q are the pre- and post-conditions, and C denotes the program being executed. In seL4, the abstract behavior of kernel operations is modeled in a monadic style that yields imperative-like programs; the corresponding Hoare rules capture the semantics of these monadic computations.
[…]
To reduce the burden of manually constructing proofs, developers typically prepare a collection of broadly applicable tactics and lemmas, dedicated to software verification. For example, the wp tactic, short for weakest precondition, generates the verification conditions associated with a given Hoare triple. During this process, wp systematically applies a suite of previously proved lemmas, each of which encodes a weakest-precondition characterization for a specific construct or monadic computation. After a routine use of simp to unfold cap_swap, human experts can then easily invoke the pre-defined tactic wp together with the lemma set_cap_valid_objs.
In this example, the target theorem is discharged by combining the tactics wp and simp. The simplification step performs splitting, but the usual splitting rule for if is disabled to avoid unnecessary case distinction. In the proof script, the symbol | indicates that the tactic may apply either branch (or both), while the trailing + instructs it to repeat this process until no further progress is possible. This interleaving enables wp to generate obligations that simp can reduce, while simp reshapes the goals to facilitate subsequent wp reasoning. In this case, the combined step eliminates the goal entirely, and the proof concludes with done.
To assess LLMs in generating this proof automatically, we experiment with two state-of-the-art proprietary models, i.e., Gemini 3 Pro (google2024gemini, ) and GPT-5.1 (openai2024gpt4techreport, ); their results are shown on the right panel of the figure. While both deep-thinking models successfully recognize the seL4 proof context, they are unable to apply the domain-specific tactic wp correctly. They also hallucinate lemmas that do not exist in the library (e.g., hoare_seq_ext). We further attempted several rounds of iterative repair by feeding back the ITP’s error messages, but the regenerated proofs continued to fail verification. An analysis of the models’ reasoning traces shows that they frequently misinterpret the functionality and performance characteristics of the wp tactic and related lemmas (such as set_cap_valid_objs).
There are two potential approaches to addressing this issue. The first is to use retrieval-augmented generation (yang2023leandojo, ; thompson2024rango, ). In this paradigm, given a target theorem, a retrieval module searches for relevant or structurally similar theorems and proofs, which are then supplied to the LLM as contextual guidance. However, this strategy depends on both a high-quality proof corpus and an effective retrieval model: requirements that are difficult to satisfy in our setting. For example, in seL4, when developers introduce a new module (such as SysInit, as shown in our evaluation) together with its accompanying proofs, comparable proofs are often absent from existing libraries. Moreover, even when analogous theorems do exist, neither model-agnostic nor model-dependent retrieval techniques reliably identify them within a large verification codebase (berlot2025llm, ; sesterhenn2025compute, ). In such cases, weak retrieval can degrade, rather than improve, the quality of generated proofs.
Another approach is to fine-tune LLMs on collected proof data. This method is considerably easier to implement and has already demonstrated promising results in formal mathematical proving. However, fine-tuning a model to directly generate complete proofs for software verification remains highly challenging.
First, even in seL4, the available corpus consists of only about 20K theorems, which severely limits the effectiveness of fine-tuning. Second, unlike formal mathematics, most proofs in systems like seL4 are procedural: as our example illustrates, their true semantics reside implicitly in the theorem prover’s internal proof states rather than in the proof scripts themselves. As a result, LLMs struggle to acquire the necessary reasoning signal from surface-level proof text alone, making this approach impractical for now.
3 Methodology
We propose a neuro-symbolic approach for proof step search on systems-level theorems. The overall procedure is illustrated in Figure 2. Starting from the root node, which represents the initial proof state, the search alternates between proposing new steps and deriving new proof states by executing those steps within the ITP.
To construct the next set of candidate proof steps for the current proof state, we first prompt a fine-tuned LLM to generate potential steps, and then apply a rule-based revision to refine or fix any imprecise suggestions. After executing these steps and obtaining a collection of new proof states, we introduce a hybrid mechanism to identify the most promising ones for further exploration. Specifically, we apply QuickCheck (claessen2000quickcheck, ) and Nitpick (blanchette2010nitpick, ), which serve as counterexample generators that help detect false conjectures or invalid intermediate proof states in Isabelle/HOL, to discard any proof state that contains a counterexample. We then use the LLM’s prediction score (cumulative perplexity) to rank the remaining states and select the most promising one. Finally, if the search for the proof fails, we invoke an automated tool, i.e., Sledgehammer, to attempt to complete the proof. Putting all the above together, we establish the framework in Algorithm 1.
Algorithm 1 Proof Search Framework
1:initial state $s_{0}$, neural generator $\mathcal{G}$, symbolic checker $\mathcal{C}$, max iterations $I_{\max}$
2:initialize tree $T$ with root containing $s_{0}$
3:for each iteration $i\in[1,I_{\max}]$ do
4:  $n\leftarrow$ highest_score_unexplored($T$)
5:  succStates $\leftarrow[]$; failStates $\leftarrow[]$
6:  candidates $\leftarrow\mathcal{G}(n.state)$ $\triangleright$ step generation
7:  for each step $\delta\in$ candidates do
8:   $s\leftarrow$ apply($\delta$, $n.state$) $\triangleright$ ITP execute
9:   if $s$ is error state then
[…]
11:   else
[…]
13:  for each $(s,\delta,\log p_{\delta})$ in Revise(failStates) do
14:   $s\leftarrow$ apply($\delta$, $n.state$)
[…]
16:  for each $(s,\delta,\log p_{\delta})$ in succStates do
17:   if s has no subgoals then
18:     return proof script $\triangleright$ proof completed
19:   if $\mathcal{C}(s)$ detects duplicate or counterexample then
20:     continue $\triangleright$ state filtering
21:   score $\leftarrow$ combine($n.score$, $\log p_{\delta}$) $\triangleright$ scoring
22:   insert_node($T$, $(s,score)$)
23:return FAIL
3.1 Proof Step Generation & Revision
3.1.1 Proof Step Generation
Our approach begins by generating candidate proof steps through prompting a fine-tuned LLM. Given the current proof state, which encapsulates the available hypotheses and the target goal, we query the model to propose the next proof step using a simple instruction template:
### Given the following Isabelle proof state,
suggest the next proof step.
### Input:
<current_proof_state>
### Response:
<next_proof_step>
The rationale for focusing on the current proof state is twofold. First, automated theorem proving can be cast as a sequential decision-making problem, and thus can be modeled as a Markov decision process, in which a standard result stipulates that pure, history-independent polices (i.e., the next step depends solely on the present state) suffice for the reachability objective, i.e., to reach the “proof done” state. This is also a common practice in the literature of proof search, which turns out to be critical for efficiency (wu2021tacticzero, ; huang2025leanprogress, ; wu2020neural, ). Second, the structure and vocabulary of the proof state provide strong cues regarding which tactics are applicable and which premises are likely to be relevant (gauthier2021tactictoe, ; piotrowski2023machine, ). Additionally, we focus on predicting a single step at a time. This is because, although multi-step prediction is possible in principle, the complexity of real proofs and the limited amount of available data make such predictions considerably more challenging; empirically, one-step prediction yields more reliable results.
To fine-tune the model to predict such next steps, we construct a dataset of state–step pairs extracted from existing theorems and their complete proofs. Each proof is decomposed into individual proof steps and replayed in the ITP, allowing us to record the intermediate proof state preceding every step. Using these aligned pairs, we fine-tune the LLM to generate outputs that are not only syntactically correct but also semantically sensible. Finally, since proof search demands efficient generation, we employ relatively small LLMs (e.g., 1.7B or 7B size), which are nevertheless sufficient for our specific tasks.
3.1.2 Proof Step Revision
Writing syntactically and semantically valid proof steps in an ITP (such as Isabelle) is challenging, particularly in system verification tasks that depend on numerous domain-specific lemmas and tactics. Consequently, even a well-trained LLM may generate steps that either refer to premises or tactics absent from the current proof context, or cannot be applied to the current goals due to syntactic errors or lack of progress. Such failures substantially degrade the effectiveness and efficiency of the proof search.
Despite their incorrectness, these failed attempts often contain useful information. Continuing our running example, an LLM might identify relevant premises cap_swap_def but pair them with an inapplicable tactic by; or it may propose a promising tactic wp while selecting premises hoare_seq_ext that do not precisely fit the context.
To fully exploit these incorrect but reasonable responses from LLMs, we employ a best-effort repair procedure. Whenever the ITP rejects an LLM-generated step, we record it for post-processing and re-evaluate the revised version. After examining all proposed steps for a given proof state, we systematically transform the failed ones in two ways: (1) tactic repair: we extract the premises referenced by misused tactics and recombine them with a curated set of frequently used tactics; and (2) premise repair: we search the proof context for similar premises and substitute them where appropriate. By synthesizing additional candidate steps, these symbolic revisions help reveal prover-accepted steps, thereby strengthening the overall proof search.
Algorithm 2 Proof-step revision and execution
1:Failed proof state set failStates
2:revisedSteps $\leftarrow[]$
3:premiseSet $\leftarrow$ established from trainingCorpora
4:tacticSet $\leftarrow$ established from trainingCorpora
5:for each $(s,\delta,\log p_{\delta})\in$ failStats do
6:  if $s$.error is w.r.t. tactic then $\triangleright$ tactic repair
7:   $p$ $\leftarrow$ extract referred premises in $\delta$
8:   revisedSteps.add(combine(tacticSet, p))
9:  if $s$.error is w.r.t. premise then $\triangleright$ premise repair
10:   $t/p\leftarrow$ extract referred tactic/premises in $\delta$
11:   $p^{\prime}\leftarrow$ retrieve similar premises from premiseSet
12:   revisedSteps.add(combine($t$, $p^{\prime}$))
13:return revisedSteps
For the tactic repair, we first extract all tactics from the training corpus. Given a proof step rejected by the ITP, we attempt to rewrite it by identifying the premises referenced in the failed step and recombining them with the pre-built tactic set. As to the premise repair, which explicitly targets proof steps that fail with an “undefined fact” error, we attempt to correct every LLM-generated undefined premise using an available version from the proof context. Concretely, we compare the undefined fact with all candidate premises using edit distance and select the closest matches (e.g., the top three). The failed step is then revised by substituting the matched premises for the undefined ones.
The procedure of proof-step revision is summarized in Algorithm 2. However, the revision can produce a large number of candidate steps. To alleviate this, we impose constraints to keep the search space manageable. For tactic repair, we limit the tactic set (e.g., to the 12 most frequently used tactics). For premise repair, we pre-select a compact subset of premises from the library. In particular, we use MePo (a heuristic module in Isabelle’s Sledgehammer that ranks and selects premises relevant to a given goal) as a lightweight relevance filter to retrieve the 128 facts most related to the current subgoal.
3.2 Proof State Filtering & Ranking
3.2.1 Proof State Filtering
To alleviate state explosion in tree-search approaches, we reorganize the accumulated proof states before further exploration. Particularly, after generating new proof states by executing the newly produced proof steps, we first invoke two complementary symbolic tools, QuickCheck (claessen2000quickcheck, ) and Nitpick (blanchette2010nitpick, ), to detect potential counterexamples in the candidate states. QuickCheck performs random property-based testing on executable parts of the proof state, whereas Nitpick translates higher-order formulas into finite relational models and searches for counterexamples using Kodkod, which is a SAT-based first-order relational model finder.
Additionally, we adapt the tool SolveDirect (SolveDirect2025, ) to detect potentially duplicate proof states. Initially, SolveDirect determines whether a newly stated theorem can be solved directly using an existing one. In our setting, we formalize two proof states as theorems and apply SolveDirect in both directions to determine whether they are semantically equivalent.
We use a basic theorem in seL4 as an example to illustrate proof state filtering. Consider the “signed overflow” theorem, whose formal statement is as follows.
theorem sofl_test:
sint x + sint y = sint (x + y) $\Longleftrightarrow$
drop_bit (size x - 1)
((x + y XOR x) AND (x + y XOR y)) = 0
for x y :: a::len word
The theorem claims that for two signed machine words, the value of their machine-level sum matches the sum of the mathematical values they represent. This property guarantees the correctness of the machine-word definitions on which the low-level kernel specifications depend. To prove this theorem, the LLM proposes a total of 8,445 candidate proof steps, rendering exhaustive exploration prohibitively time-consuming. Nevertheless, our filtering reveals that 44.2% of these steps lead to duplicate proof states, and 52.3% of the remaining states are unprovable due to existing counterexamples. Thus, the state-filtering process effectively eliminates unproductive effort and conserves the search budget, enabling the proof to be found much earlier.
3.2.2 Proof State Ranking
To guide the proof search efficiently, we rank all proof states and always expand the most promising ones first. To score a given proof state, we rely on the log probabilities predicted by the fine-tuned LLM. The LLM is trained to predict the next proof step; its log probability directly reflects the model’s confidence in applying that step to the current proof state.
Consider a candidate proof state $s_{L}$, obtained by applying a sequence of proof steps $(a_{1},\dots,a_{L-1})$. We define its cumulative log-probability score as
$\text{score}(s_{L})=\frac{\sum_{t=0}^{L-1}\log p(a_{t}\mid s_{t})}{L^{\alpha}},$
where $s_{1},\dots,s_{L-1}$ are the intermediate states produced by the corresponding steps. The normalization term $L^{\alpha}$ mitigates the inherent bias against longer proof sequences, with the exponent $\alpha$ specifying the degree of mitigation xin2025bfs , which we set to 1 by default. After computing the score for each proof state, we select the top-k states for further expansion where $k$ is a hyperparameter.
3.3 Hammer Integration
Hammers are tools designed to discharge proof goals in ITPs automatically. In Isabelle, the primary hammer is Sledgehammer bohme2010sledgehammer , which applies ATPs and SMT solvers (e.g., Z3, CVC5, E, SPASS, Vampire, etc) to the current goal and synthesizes corresponding Isabelle proofs. When invoked, it first retrieves a fixed number of potentially relevant premises that may aid in solving the goal. These premises, together with the goal, are translated into a form suitable for backend ATP and SMT solvers. If a backend solver succeeds, Sledgehammer attempts to reconstruct a corresponding Isabelle proof step, and, when this reconstruction succeeds, a complete proof for the current goal is returned to the prover.
Although Sledgehammer’s performance is far from satisfactory for our tasks, it plays a mutually beneficial role in the tree search process. In particular, while tree search may fail to solve the original problem within a limited search budget, it often reduces the problem to a simpler form that Sledgehammer can solve, even though it cannot solve the initial goal on its own. This makes it natural to integrate the hammer into our framework. Because invoking the hammer at every search step is prohibitively expensive, we only call it when tree search fails. Specifically, upon search failure, we rank all proof states in the currently expanded tree, select a small subset with the highest scores, and pass them to Sledgehammer for a final attempt at solving the goal.
4 Evaluation
This section presents our experimental results. We use seL4 as a system-level evaluation testbed to demonstrate both the effectiveness and efficiency of our proposed method. In addition, we introduce four additional repository-level benchmarks to further assess its generalizability. Specifically, our study aims to address the following research questions (RQs):
  • RQ1 – Effectiveness: Does our proposed method outperform existing techniques in generating proofs for seL4?
  • RQ2 – Efficiency: How much human effort in proving seL4 can be saved using our proposed method?
  • RQ3 – Generalizability: Can our method perform well on the verification of other software projects?
  • RQ4 – Ablation: What is the contribution of each component within our framework?
All evaluations were carried out within a dedicated Docker environment on a Linux server with an AMD EPYC 9654 96-core processor and six high-end GPUs.
4.1 Isabelle REPL
To enable efficient interaction with Isabelle from Python, we implement a new Isabelle REPL (Read-Eval-Print Loop), [^2] Isabelle REPL is an interactive command-line interface for the Isabelle theorem prover, designed to support incremental, exploratory proof development, which is similar in spirit to the REPLs found in programming languages such as Python, but adapted to Isabelle’s document-oriented proof model. whose logic builds on scala-isabelle unruh2025scalaisabelle and Py4J (nishant2023py4j, ). The REPL provides a Gateway Server that exposes Isabelle components to Python clients through a unified API. This design enables programmatic interaction with Isabelle, allowing users to parse Isabelle theories, extract run-time context information, and execute proof steps incrementally.
Compared to the previously developed Isabelle REPL, PISA jiang2021lisa , our REPL supports more recent Isabelle versions and provides extended ML-level capabilities. A wide range of Isabelle’s proof-automation tools, including Sledgehammer, Nitpick, Quickcheck, and other integrated methods, can be accessed directly through our interface. It also improves the extraction of information from proof contexts, such as variables and assumptions in proof goals, dependencies of proved theorems, and facts selected by Sledgehammer.
In addition, our REPL provides improved management of the Isabelle process, top-level state, and theories. It allows users to safely impose strict time limits on Isabelle processes, as well as to clone, restore, and switch between different proof states. Theory management is also strengthened: in large verification projects, complex theory dependencies can heavily degrade compilation efficiency. For example, in the CRefine theory of seL4, a single proof step takes more than one minute to process. To address this, we implement a caching mechanism that avoids redundant execution of unrelated theorems and proofs, significantly reducing unnecessary recomputation.
4.2 Experimental Setup
Dataset. We use the FVELER dataset (lin2024fvel, ), which comprises 29,125 theorems from seL4. The dataset is split into four subsets, i.e., training, validation, test, and test-hard, containing 26,081, 1,115, 1,077, and 852 theorems, respectively. Following the FVEL criteria, the training, validation, and test sets are randomly partitioned. In contrast, the test-hard set consists of theorems drawn from three specific and independent sessions, i.e., SysInit, SysInitExamples, and LibTest, that do not appear in any of the other splits. Note that sessions are the way Isabelle organizes theory files, which resemble the relationship between code files and libraries in other programming languages. A theory can import other theories, and a session can depend on other sessions by declaring dependencies in the Isabelle ROOT file. These sessions are typically selected based on the depth of their session-level dependency graphs, and the theorems they contain exhibit more complex dependencies, making them considerably more challenging to prove.
In our experiments, we use the training set as the corpus for fine-tuning LLMs, and adopt validation, test, and test-hard sets to evaluate our proposed framework. From the training set, we further extract all proof state–step pairs, yielding a total of 181,887 pairs. On average, each theorem contributes to about seven pairs, substantially increasing the amount of supervised data available for model training.
Training. We fine-tuned two small models, Qwen3-1.7B (yang2025qwen3, ) and Mistral-7B (jiang2023mistral, ), to predict and score each proof step. We carried out the full-parameter supervised fine-tuning (SFT) using the Llama-Factory framework (zheng2024llamafactory, ). We use DeepSpeed ZeRO-2 optimization to enable memory-efficient distributed training. The training used an effective global batch size of 16 and ran 3 epochs using bfloat16 mixed precision. The learning rate was set to 1e-5 and followed a cosine decay schedule with a warmup ratio of 0.1.
Proving. For the LLM inference, we adopt a high temperature of 1.0 and a top-p value of 0.95 to explore the proof space sufficiently. The maximum generation length is fixed at 2,048 tokens. In each search iteration, at most five proof states are selected; for each proof state, the model is prompted to generate 128 candidate proof steps. In our framework, we disable LLMs’ internal thinking to improve inference efficiency.
When proof search fails, we select the 16 highest-scoring proof states and invoke Sledgehammer for assistance. We run Sledgehammer with its default premise-selection strategy: it first selects 2,048 most relevant premises for the current proof goal using MeSh, which combines MePo meng2009lightweight and MaSh kuhlwein2013mash to gauge relevance between the proof goal and available facts in the context. With these retrieved results, Sledgehammer then attempts to discharge the current subgoal using the built-in SMT/ATP solvers Z3, CVC5, E, SPASS, and Vampire, with a 60s time limit.
Baseline. We select four neural or symbolic proof-generation methods as baselines. Selene prompts GPT-4o with the target problem and several examples to generate a candidate proof. FVEL follows a similar paradigm but also fine-tunes a Mistral 7B-Instruct model on the training data to produce complete proofs, rather than using in-context learning. We also include two symbolic methods, Auto and Sledgehammer, which operate purely symbolically to prove the target theorem. Here, Auto is a collection of common symbolic automatic proving methods in Isabelle, such as by simp, by auto, and by fastforce. The time limit for all methods is set to 120 minutes, as we observed no noticeable improvement when increasing it further.
4.3 Experimental Results
4.3.1 RQ1: Effectiveness Comparison
Table 1: Proof success rates (%) across validation (Val), test (Test), and test-hard (TsHd) splits. Compared to existing neural and symbolic provers, our framework delivers substantial improvements: Qwen3-1.7B achieves 70.4% overall success, and Mistral-7B reaches 77.6%, exceeding the strongest baseline by 30.0% and 37.3%, respectively.
Method Val Test TsHd Total
Neural Selene 6.1 7.0 3.3 5.6
FVEL 8.9 9.5 4.5 7.8
Symb. Auto 4.9 6.7 6.1 5.9
Hammer 40.5 39.5 40.9 40.3
Ours Qwen3 73.6 74.9 61.1 70.4
[…]
Mistral 79.8 89.0 69.8 77.6
[…]
Table 1 summarizes the performance of our approach compared to both neural and symbolic baselines. Among neural methods, Selene achieved 156 successful proofs, while FVEL improved this to 219. As for symbolic baselines, Auto proved 164 theorems, whereas Sledgehammer reached 1,124. In contrast, our framework substantially outperforms all these baselines across different splits. Using the fine-tuned Mistral-7B model, our approach produces a total of 2,167 feasible proofs, of which 788 are on the validation set, 811 are on the test set, and 568 are on the test-hard set. This accounts for 77.6% of all still valid theorems across the three evaluated sets and proves 37.3% more theorems than the best baseline, Sledgehammer. It is also encouraging that, on a relatively small model, Qwen3-1.7B, our approach manages to complete the proof of 1,965 theorems, which amounts to 70.4% of the total theorems. This demonstrates that in a real environment where new proofs need to be crafted, a deployed small model can already help with the development of verification.
We further evaluate the performance of our automated proving approach in two aspects. First, prior work has struggled with generating nontrivial proofs. Since the proofs in seL4 are largely procedural, we use the proof length as a proxy for proving difficulty. We compute the proof success rate across different proof lengths and present the results in Figure 3. The results show that our approach is more capable of proving longer theorems than previous baselines. Although the success rate naturally decreases as proof length increases, it remains around 20% even for the 393 theorems whose proofs exceed 10 lines. We leave it as future work to investigate more effective strategies for constructing longer proofs.
Figure 3: Proof success rate across ground-truth proof lengths. Our approach consistently outperforms existing methods. In particular, the success rate decreases as proofs grow longer, but our approach still maintains non-negligible performance even on proofs exceeding 10 lines.
[…]
Response to RQ1 Our approach substantially outperforms baselines, producing 37.3% more proofs than the (symbolic) ATP tool Sledgehammer and 69.8% more proofs than the state-of-the-art neural method. It remains reliably effective across different proof difficulty levels and session categories. Further, it demonstrates good generalizability, successfully discharging 67.6% of the theorems from unseen sessions during training.
Figure 4: Proof success rates across the six seL4 session categories. Our approach yields stronger performance than all baselines in all sessions. Notably, it achieves high accuracy even on the challenging SysInitGroup category, which is entirely unseen during training.
Table 2: Coverage results of automatically generated proof lines. Our method attains the highest coverage, proving theorems whose existing proofs amount to over one-third of all proof lines and substantially outperforming existing baselines in reducing manual proof-script development effort.
Method # of proof lines % of coverage
Selene 197 1.1
FVEL 271 1.6
Auto 204 1.2
Hammer 2,581 15.0
Ours 6,235 36.2
Figure 5: Proof completion rate under varying proportions of expert-provided proof lines. Our method is increasingly effective as more human-written lines are supplied. Overall, the method reduces human effort by 71.1% on average.
4.3.2 RQ2: Proving Efficiency in Saving Human Efforts
To evaluate how effectively our approach reduces manual effort in developing proof scripts, we adopt a standard metric: the coverage of the entire library that can be automatically generated matichuk2015cost . As summarized in Table 2, our method successfully proves theorems whose corresponding proofs account for 6,235 lines in the seL4 proof corpus, achieving 36.2% coverage across all evaluated theorems. Under this metric, our approach yields more than a one-fold improvement over existing baseline methods.
We further situate our method within an AI–human collaborative pipeline to evaluate its efficiency in reducing human effort. Specifically, for each theorem, we provide a portion of the ground-truth proof lines (simulating human expert-written proof) and ask our approach to complete the remaining steps. Figure 5 shows the proof completion rate curve w.r.t. the proportion of ground-truth proof.
As expected, supplying more expert-written proof lines enables our method to complete more theorems. Ultimately, including cases solved entirely automatically, our approach succeeds in automatically completing at least 10% export effort for 86.0% of the theorems. We then calculate the average effort saving (AES) of our approach:
$\text{AES}=p_{1}\sigma_{1}+\sum_{i=2}^{k}(p_{k}-p_{k-1})\sigma_{k},$
where $(\sigma_{1},\dots,\sigma_{k})$ are the ratios of expert-written proof lines in ascending order, with the corresponding proof completion rates are $(p_{1},\dots,p_{k})$. Therefore, under this AI–human collaboration setting, the automated proof completion reduces human effort by an average of 71.1%.
We also record the time consumption of our framework as a reference; the average time to generate a successful proof is 139.1 minutes. However, it is worth noting that this average is skewed upward by a small number of theorems that take extremely long to prove. In our experiment, we find that 58.4% of the theorems are proved within 10 minutes, 73.4% within 30 minutes, and 80.8% within 2 hours. Therefore, the majority of proofs are obtained within a relatively short time horizon, indicating that our framework is generally highly efficient despite a few outliers with long proof times.
Response to RQ2 Our approach alleviates much of the burden of interactive theorem proving, automatically completing proofs that correspond to 36.2% of the proof lines in the evaluated corpus and reducing expert effort by approximately 71.1% in an AI–human collaboration setting. Moreover, it is able to synthesize the majority of proofs within a practical time budget.
4.3.3 RQ3: Generalizability to Other Projects
A major concern with training-based methods is their generazability, namely, whether the approach is only applicable to a specific domain due to potential data leakage or overfitting.
To investigate possible leakage, given that the seL4 proofs are publicly available and that the pre-trained model may have already incorporated some of the knowledge, we compute the similarity between the ground-truth proofs and the proofs generated by our approach. We adopt sequence similarity (lcvenshtcin1966binary, ) and Jaccard similarity (jaccard1901etude, ) as the evaluation metrics. The results are shown in Figure 6. They indicate that both similarity scores remain low except for one- or two-line proofs, which is not surprising since very short proofs tend to be identical. We further examined these short proofs and found that only 16.7% are identical, providing strong evidence that data leakage in our fine-tuned model is unlikely.
Figure 6: Sequence and Jaccard similarity between generated and ground-truth proofs. (Note that sequence similarity measures how alike two ordered sequences are based on their matching elements and positions, while Jaccard similarity measures how similar two sets are by comparing the proportion of shared elements to the total distinct elements.) Both similarities stay below the thresholds of indicating potential data leakage across different proof lengths.
We further explore the generalizability of our approach and the fine-tuned LLM. We select three projects from the Archive of Formal Proofs (AFP [^3] http://afp.sf.org), i.e., X86 Semantics X86_Semantics-AFP , IEEE Floating Point IEEE_Floating_Point-AFP , and SATSolverVerification SATSolverVerification-AFP , to directly apply proof search on the theorems therein. In addition, we include the code verification benchmark Code2Inv and translate its loop invariants into Isabelle theorems.
The performance of our approach and baselines is shown in Figure 7. Because the X86 semantics benchmark is more relevant to the seL4 project, the fine-tuned LLM performs particularly well on it, yielding a 32.5% higher success rate than the baselines. The other three benchmarks are hammer-friendly because of their SMT-oriented nature. Even in this setting, our framework demonstrates superior effectiveness, achieving an average relative improvement of 36.9%.
Figure 7: Performance of our proof search framework on additional benchmarks. The $x$-axis lists the benchmarks, where entries in brackets denote the number of theorems included in each benchmark. Our approach consistently achieves higher proof success rates across all benchmarks, demonstrating strong generalizability and robustness.
Response to RQ3 The knowledge acquired by the fine-tuned LLM, together with our proof search framework, generalizes well across domains. In particular, the approach can be effectively leveraged to assist proof construction in a wide range of formal verification tasks beyond the seL4 setting.
4.3.4 RQ4: Ablation study
Since our approach integrates several components, namely LLMs, hammer, as well as proof-step revision and proof-state filtering mechanisms, we conduct ablation studies to evaluate the contribution of each part of the framework. To assess the impact of our fine-tuned model, we replace it with off-the-shelf LLMs such as GPT-4o or DeepSeek. Due to budget constraints (running tree search for a single problem costs approximately \$4.1 when using the DeepSeek or GPT-4o APIs), we randomly sample 200 theorems from the benchmark and use the DeepSeek API to propose proof steps.
In addition, inspired by Rango (thompson2024rango, ), we implement a retrieval-augmented generation (RAG) mechanism. The retrieval module fetches both similar proof steps and potentially relevant premises. Similar steps are retrieved from the training set based on the similarity between their associated proof states and the current proof state, where these states are tokenized into sequences of identifiers, and relevance is measured using BM25 (robertson2009probabilistic, ). For premises retrieval, we continue to rely on Sledgehammer’s MePo heuristic for efficiency. The final prompt for LLMs includes the retrieved steps and premises along with their statements, and we restrict the total lengths of steps and premises to 1024 and 512 tokens, respectively.
The results for all method variations are reported in Table 3. Overall, although adding tree search and RAG steadily improves DeepSeek’s performance, these variants still fall significantly short of our full framework. This confirms that the combination of the fine-tuned model, structured search, and hammer integration is essential for achieving high proof success rates, and highlights the importance of each major component in our design.
To evaluate the impact of the step revision mechanism, we analyze the theorems that become difficult to prove when the revision is disabled. We observe that revision is most beneficial when the LLM fails to generate enough effective proof steps. In particular, for the most challenging 77 theorems where the model initially proposed no valid steps and therefore failed immediately, the revision procedure is highly effective, enabling 24.7% of these theorems to be proved. For a broader set of 220 theorems where the LLM proposed fewer than 5% valid steps, the revision mechanism still provides substantial gains, allowing 11.8% of them to succeed.
Regarding the necessity of proof state filtering, particularly the counterexample detection, our experiments show that incorporating this component allows the system to identify an average of 1.3 counterexamples per evaluated theorem. For the subset of search traces that contain counterexamples, the average number rises to 11.1. These results indicate that counterexample detection is valuable for identifying unprovable states early and preventing unproductive search efforts, although it does not directly increase the overall proof success rate.
Response to RQ4 Each of the symbolic and neural components in our framework plays an essential role, and their synergy substantially strengthens the overall theorem-proving capability for software verification.
Table 3: Proof success rate (%) on several variants of our framework. The proof success rate increases progressively as tree search, retrieval augmentation, and hammer are incorporated, indicating the necessity of each component. The comparison highlights that our full approach, which replaces the off-the-shelf model with a fine-tuned LLM, achieves substantially stronger performance.
Method Variation Rate
DeepSeek (DpSk) v3.2 5.5
DpSk + Tree Search 12.5
DpSk + RAG + Tree Search 33.0
DpSk + RAG + Tree Search + Hammer 52.0
Ours (Trained LLM + Tree Search + Hammer) 70.0
5 Related Work
5.1 Software System Verification via ITP
Formal methods have successfully verified critical software systems including operating system kernels, hypervisors, file systems and compilers. Interactive theorem proving (ITP) provides strong formal assurance by allowing engineers to encode specifications and build proofs in ITP systems such as Rocq or Isabelle/HOL. Notable examples include CertiKOS (gu2016certikos, ), which employs a compositional refinement framework to verify a layered OS kernel; CompCert leroy2016compcert , an optimizing C compiler whose translation passes are all verified to preserve semantics; FSCQ (fscq, ), a crash-safe file system proven correct under all failure scenarios using Crash Hoare Logic; seKVM hypervisor (sekvm, ), a reduced KVM core with enforced confidentiality and integrity of guest memory; seL4 klein2014comprehensive , an OS microkernel which remains the most comprehensive and influential milestone for formal system verification. These efforts demonstrate that ITP enables high-assurance proofs for critical systems, albeit at a significant engineering cost.
5.2 Automated Theorem Proving
Automated theorem proving (ATP) has long served as a benchmark for machine reasoning. Classical ATP systems rely on symbolic methods, such as resolution-based first-order provers (e.g., Vampire (riazanov2002design, ), E (schulz2019faster, )) and SMT solvers (e.g., Z3 (de2008z3, ), CVC5 (barbosa2022cvc5, )) with built-in theories. These solvers are also integrated into interactive theorem provers via hammers (e.g., Sledgehammer (bohme2010sledgehammer, ) in Isabelle), enabling automated discharge of proof goals (czajka2018hammer, ). Related to ATP, several works (e.g., Ironclad Apps (hawblitzel2014ironclad, ) and Verve (yang2011safe, )) leverage verification-aware languages (e.g., Dafny (leino2010dafny, ) and Boogie (barnett2005boogie, )) to prove memory and control-flow safety of system components, including kernels and device drivers.
In recent years, machine learning has been increasingly applied to ATP. Early efforts focused on premise selection and tactic prediction using classical learning methods (e.g., Holophrasm (whalen2016holophrasm, ), TacticToe (gauthier2021tactictoe, ), Proverbot9001 (sanchez2020proverbot, ), Diva (first2022diversity, )), showing improved success rates in Metamath, HOL4, and Coq. More recently, large language models (LLMs) have been utilized to generate complete proofs or sketches (jiang2022draft, ; cao2025reviving, ), achieving possible superhuman results on mathematical problems (trinh2024solving, ). Prior attempts to extend these capabilities to software verification include Baldur (first2023baldur, ), PALM (lu2024proof, ), and Guided Proof Search (prasad2024guided, ). However, they remain limited to textbook-scale problems, which are insufficient to address challenges from real-world systems-level verification projects such as seL4.
Complementary to whole-proof generation, the proof-step prediction paradigm harnesses LLMs to propose the next tactic or lemma, guided by proof-state feedback. Systems such as GPT-f (polu2020generative, ), LISA (jiang2021lisa, ), HTPS (lample2022hypertree, ) and LeanDojo (yang2023leandojo, ) demonstrate the potential of this class of methods across Metamath, Isabelle and Lean. However, this line of work primarily targets singleton mathematical problems and therefore does not scale to system-level software verification. Our method fills this gap by integrating a fine-tuned LLM for proof search with symbolic techniques, specifically, proof-step revision and proof-state filtering, making large-scale verification both effective and efficient.
To further enhance reliability, hybrid approaches are proposed to incorporate symbolic provers into LLM pipelines. A series of works, e.g., Thor (jiang2022thor, ), ProofAug (liu2025proofaug, ) and Strat2Rocq (fang2025proof, ), exemplify this by coupling LLM-generated reasoning with ATPs or hammers. Yet, ATP-based verification fundamentally trades expressiveness for automation, limiting scalability in system-level contexts. In contrast, our method integrates LLMs with an ITP workflow, invoking ATPs only in the final stage to discharge the remaining proof obligations.
6 Limitations and Future Work
Despite strong empirical performance, our framework has several limitations that suggest directions for future research.
Computational cost. The overall tree-search procedure remains computationally expensive. Each iteration requires invoking Isabelle for proof-step execution and proof-state checking, which results in substantial runtime overhead. Future work includes developing efficient state caching, parallel expansion, and cost-aware search policies to minimize unnecessary exploration and improve throughput.
Performance degradation for longer proofs. While our method successfully constructs longer proofs than prior methods, its performance indeed degrades as proof length increases. The current tree-search procedure adopts one-step prediction and bounded search, making it challenging to uncover deeper reasoning chains. Future improvements could involve hierarchical planning, subgoal decomposition, or reinforcement learning from bootstrapped trajectories.
Rigid rule-driven step revision. The proof step revision modules heavily rely on heuristic tactic-premise recombination and lightweight premise retrieval. These mechanisms, though helpful, are imperfect and may introduce misleading candidates. Future improvements may include learning the relevance of premises directly from the proof state or integrating stronger semantic filters.
7 Conclusion
In this paper, we have presented a neuro-symbolic proof-generation framework that automates proof search for large-scale software verification projects in interactive theorem provers. Instead of attempting whole-proof synthesis, our approach performs a best-first tree search over fine-grained proof states, repeatedly querying a lightweight, fine-tuned LLM for the next proof step. This paradigm not only substantially improves data efficiency, but also enables a seamless integration of symbolic reasoning tools with LLM-based step generation and scoring, forming an effective and efficient pipeline for proof-step construction as well as proof-state filtering and ranking. The experiments on the seL4 verification corpus show our approach significantly outperform prior baselines. Notably, our method succeeds on a considerably higher proportion of multi-step, nontrivial theorems, addressing a long-standing limitation of LLM-based provers. Additional experiments on AFP projects and translated program-verification benchmarks further showcase strong cross-project generalizability.
Our work provides compelling evidence that proof-step–centric neuro-symbolic search can serve as a practical path toward scalable and automated software verification at the systems level, e.g. operating systems kernels, network architecture, and databases. Although there is still a long way to go, we are making promising progress towards developing AI-enabled, rigorous engineering methods for real-world computer systems.
Availability
The tool and all experimental data are available at the following private link, [^4] https://figshare.com/s/da4d1d995a9aa64eeadf?file=60276236 and a prebuilt Docker image containing the seL4 verification environment and all required dependencies will also be released.